# Cancer-Myth: Evaluating Large Language Models on Patient Questions with False Presuppositions

### v1

@@ Abstract
Cancer patients are increasingly turning to large language models (LLMs) as a new form of internet search for medical information, making it critical to assess how well these models handle complex, personalized questions. However, current medical benchmarks focus on medical exams or consumer-searched questions and do not evaluate LLMs on real patient questions with detailed clinical contexts. In this paper, we first evaluate LLMs on cancer-related questions drawn from real patients, reviewed by three hematology oncology physicians. While responses are generally accurate, with GPT-4-Turbo scoring 4.13 out of 5, the models frequently fail to recognize or address false presuppositions in the questions—posing risks to safe medical decision-making. To study this limitation systematically, we introduce Cancer-Myth, an expert-verified adversarial dataset of 585 cancer-related questions with false presuppositions. On this benchmark, no frontier LLM—including GPT-4o, Gemini-1.5-Pro, and Claude-3.5-Sonnet—corrects these false presuppositions more than $30\%$ of the time. Even advanced medical agentic methods do not prevent LLMs from ignoring false presuppositions. These findings expose a critical gap in the clinical reliability of LLMs and underscore the need for more robust safeguards in medical AI systems.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive performance on various medical benchmarks, primarily focusing on medical exam questions or consumer-searched queries, such as MedQA (Jin et al., 2020) and HealthSearchQA (Singhal et al., 2023), demonstrating their ability to recall and apply medical knowledge. As a result, LLMs are increasingly utilized as real-life medical advisors. A recent survey shows that 32.6% of patients now ask LLMs for advice (Aydin et al., 2024), particularly for severe diseases like cancer, where medical resources are limited (Bourgeois et al., 2024; Belong.Life, 2024). However, the real-world patient inquiries differ significantly from traditional medical benchmarks, as they often contain patient details and sometimes false presuppositions (Kaplan, 1978), i.e., a misconception that a patient believes to be shared knowledge between themselves and the clinician when asking a question. To effectively respond to real-world patient inquiries, an LLM should (1) provide factual and helpful answers and (2) identify and clarify any misconceptions in the patient’s question. Both aspects are essential for ensuring patient safety and effective medical communication.
@@
To systematically study this issue, we compile a collection of 994 common cancer myths and develop an adversarial Cancer-Myth dataset of 585 examples, designed to evaluate LLM and medical agent performance in handling patient questions with embedded misconceptions. We initialized the adversarial datasets with a few failure examples from our previous CancerCare study. Using an LLM generator, we created patient questions for each myth, integrating false presuppositions with complex patient details to challenge the models. The LLM responder answers these questions, while a verifier evaluates the response’s ability to address false presuppositions effectively. Responses that fail to correct the presuppositions are added to the adversarial set, while successful ones are placed in the non-adversarial set, for use in subsequent generator prompting rounds. We perform three separate runs over the entire set of myths, each targeting GPT-4o, Gemini-1.5-Pro, and Claude-3.5-Sonnet, respectively. The generated questions are categorized into 7 categories, and then finally reviewed by physicians to ensure their relevance and reliability.
Our experimental results reveal a lack of awareness in identifying and correcting patient false presuppositions during LLMs’ medical reasoning. No model can fully correct more than 30% of the false presuppositions. The state-of-the-art medical agentic system MDAgents (Kim et al., 2024), aggregating information from LLMs role-playing as doctors, is ineffective on Cancer-Myth, only correcting 2.2% of the false presuppositions. In particular, models perform worst on the Inevitable Side Effect category, where patient misconceptions, such as assuming a specific treatment will inevitably cause a certain side effect, often go unchallenged by LLMs. These findings underscore the need for improved LLM training and evaluation methods that emphasize patient-centered communication and misinformation detection.
@@
2 Physician Evaluation of LLM Responses to Real Patient Questions
@@ 2.1 CancerCare data preparation
We selected 25 representative oncology questions sourced from CancerCare website [^1] https://www.cancercare.org/questions, which provides support and resources to individuals affected by cancer (See Appendix A).
The selected question focus on treatment advice and side effect, which requires medical expertise. We confirmed these questions cannot be answered through a simple Google web search. We asked three hematology oncology physicians to evaluate these questions, featuring four responses per question: three frontier LLMs GPT-4-Turbo (OpenAI, 2023), Gemini-1.5-Pro (Google, 2023), LLaMa-3.1-405B (Meta, 2024) and the human responses from the website for comparison. The human answers are provided by licensed medical social workers, who excel in offering compassionate support and general guidance but may lack the depth of medical knowledge for complex oncological queries. Their responses average 237 words in length. To ensure the physicians remained blind to the response source, we prompted LLMs to provide responses of similar length and removed identifiers such as “As an AI chatbot”.
@@ 3 Cancer-Myth Dataset Creation
We developed a dataset to systematically evaluate LLM performance in answering questions with false presuppositions. We collected 994 cancer-related myths and used LLMs to generate adversarial examples based on these myths. These examples are then verified by hematology oncology physicians to ensure validity.
@@ 3.1 Myth collection
To construct a comprehensive collection of cancer-related myths, we began by compiling a list of 127 cancer types from the official National Cancer Institute (NCI) website [^2] https://www.cancer.gov/types. For each type, we conducted online searches to identify common myths specifically related to the treatment of that cancer. All information was collected from publicly accessible online resources. For example:
@@
Figure 4: Example question and information to correct the false presuppositions per category.
@@ 3.3 Categorization and expert validation of dataset
To effectively analyze the types of false presuppositions in patient questions, we manually reviewed a subset of 76 examples from our adversarial set, identifying six major categories of false presuppositions: Only/Standard Treatment, No Treatment, Inevitable Side Effect, Causal Misattribution, Underestimated Risk, and No Symptom, No Disease. Examples not fitting these categories are classified as Others.
We prompted GPT-4o to categorize each example, achieving an agreement rate of $\sim 90\%$ on the manually annotated subset. Therefore, GPT-4o is used to categorize the remaining examples. Figure 4 lists the questions and the corresponding information to correct the false presuppositions in each category.
To prevent any category from dominating the dataset, we employ the algorithm described in Appendix B to balance the categories before they are verified physicians, ensuring a diverse and representative benchmark. Finally, hematology oncology physicians review adversarial examples to confirm that (1) the generated patient question contain a false presupposition; (2) the generated medical knowledge is medically sound and appropriately address the false presupposition.
@@ 4.1 Dataset statistics and quality
We generated a total of 1,692 adversarial examples: 868 from GPT-4o, 393 from Gemini-1.5-Pro, 429 from Claude-3.5-Sonnet and 2 manually crafted (i.e., the initial set). After categorization and label balancing, we selected 888 examples for verification by hematology oncology physicians. They filtered these down to 585 valid data points. The final distribution of question categories and generator sources, after the annotation of physicians, is presented in Figure 5.
The physicians identified two main reasons for filtering out examples: (1) Information contradicted an ongoing event. E.g., the information says “surgery is not the only option for …”, while the question indicates the patient is already undergoing surgery, a decision made by doctors; (2) Lack of sufficient patient context. E.g., the information says “surgery is the primary treatment”, but the question does not include patient age, where age affects surgical eligibility based on performance status.
@@ 4.2 Models and metrics
We use GPT-4o as the verifier to evaluate the response of seven frontier LLMs: GPT-4o (OpenAI, 2024), GPT-4-Turbo, GPT-3.5 (OpenAI, 2023), DeepSeek-R1, DeepSeek-V3 (DeepSeek-AI et al., 2025), Claude-3.5-Sonnet (Anthropic, 2024) and Gemini-1.5-Pro (Google, 2023). We follow the same scoring system as in generation (§3.2) and output a score $s$ in $\{1,0,-1\}$. Additionally, we evaluate an adaptive multi-agent framework MDAgents (Kim et al., 2024), which dynamically assigns collaboration structures among LLMs based on the complexity of the medical task, and achieves state-of-the-art performance on 7 medical benchmarks. Since LLMs should be able to address patient misconceptions without any prompts in real patient interactions, all models are evaluated in a zero-shot setup.
Our primary evaluation metric, the Presupposition Correction Score (PCS), is the average score from the verifier. To better align with human (Appendix C) and reduce the difficulty of identifying no correction ($s=-1$) and partial correction ($s=0$), we introduce the Presupposition Correction Rate (PCR), which focuses solely on fully correct scenarios.
@@
Figure 6: (a) Gemini-1.5-Pro performs the best but no frontier LLM corrects the false presuppositions in the patient question more than 30% of the time; Multi-agent medical collaboration does not prevent LLMs from ignoring false presuppositions. (b) Adversarial data generated by Gemini-1.5-Pro causes failures in GPT-4o, but data generated by GPT-4o affects Gemini-1.5-Pro less.
@@ 4.3 Evaluation results
We observe no frontier LLM corrected more than 30% of false presuppositions in patient questions (Figure 6(a)). Among all models, Gemini-1.5-Pro achieves the highest Presupposition Correction Rate (PCR) at 27.2%, followed by Claude-3.5-Sonnet at 19.8% and GPT-4-Turbo at 15.4%. Gemini-1.5-Pro outperforms Claude-3.5-Sonnet with a p-value of $2\times 10^{-5}$, and Claude-3.5-Sonnet outperforms GPT-4o with a p-value of $4\times 10^{-3}$. Suprisingly, GPT-4o shows a low performance on Cancer-Myth, with a PCR of only 5.8%. While GPT-4-Turbo is the best model for providing factual and helpful medical responses (Figure 2), these results indicate that LLM ability to correct false presuppositions does not align with medical knowledge understanding and utilization.
@@
In our experiment, the use of the MDAgents (Kim et al., 2024) does not improve performance. Responses generated through MDAgents performed worse than standalone GPT-4o outputs, suggesting that agent-based orchestration is insufficient to address embedded misconceptions.
@@
While overall performance remains limited, Gemini-1.5-Pro achieves the highest PCS among the evaluated models. Its performance advantage primarily comes from better handling of the latter three categories—Causal Misattribution, Underestimated Risk, and No Symptom, No Disease—where it demonstrates greater capacity to detect and explain misleading assumptions. These categories often involve more subtle or technical misconceptions, such as confusing symptom origins or believing that absence of symptoms negates the need for screening. Nonetheless, even in these areas, the best model still leaves substantial room for improvement.
Figure 7: Category-wise Presupposition Correction Score for each model. All models perform poorly on No Treatment and Inevitable Side Effect
@@ 5 Related Works
LLM sycophancy
Questions containing false presuppositions have long been studied in linguistic literature (Kaplan, 1978), where the appropriate and unambiguous response is a corrective indirect response that negates the false presupposition. LLMs, however, exhibit sycophancy—a tendency to align with users’ opinions (Perez et al., 2023)—making them prone to accepting and reinforcing false presuppositions in questions. Common sycophancy mitigation strategies, such as in-context learning (Zhao, 2023), precautionary prompts (Varshney et al., 2024), and augmenting contextual knowledge from LLMs (Luo et al., 2023), are ineffective in zero-shot settings (Rrv et al., 2024), which are crucial for real-world patient-LLM interactions. Recently, Yu et al. (2023) introduced a benchmark to evaluate LLM responses to false presuppositions. In the medical domain, Cancer-Myth is the first benchmark designed to assess how LLMs handle questions containing false presuppositions.
Adversarial generation
Semantic, manual, or rule-based adversarial data generation (Jia & Liang, 2017; Rajpurkar et al., 2018) predates the emergence of LLMs, yet perturbation methods remain effective for creating challenging negatives (Zhu et al., 2022; Fu et al., 2025). Recently, automatic LLM-based synthetic adversarial generation (Bartolo et al., 2021; Fu et al., 2023) and filtering (Bras et al., 2020) have enhanced model robustness against attacks and mitigated overfitting. Dynabench (Kiela et al., 2021) combines human annotation with model inspection to generate misleading examples. To be labor-efficient, we utilize an LLM verifier with physician annotation to filter hard negative questions containing false presuppositions.

### v2

@@ Abstract
Cancer patients are increasingly turning to large language models (LLMs) for medical information, making it critical to assess how well these models handle complex, personalized questions. However, current medical benchmarks focus on medical exams or consumer-searched questions and do not evaluate LLMs on real patient questions with patient details. In this paper, we first have three hematology-oncology physicians evaluate cancer-related questions drawn from real patients. While LLM responses are generally accurate, the models frequently fail to recognize or address false presuppositions in the questions, posing risks to safe medical decision-making. To study this limitation systematically, we introduce Cancer-Myth, an expert-verified adversarial dataset of 585 cancer-related questions with false presuppositions. On this benchmark, no frontier LLM—including GPT-5, Gemini-2.5-Pro, and Claude-4-Sonnet—corrects these false presuppositions more than $43\%$ of the time. To study mitigation strategies, we further construct a 150-question Cancer-Myth-NFP set, in which physicians confirm the absence of false presuppositions. We find typical mitigation strategies, such as adding precautionary prompts with GEPA optimization, can raise accuracy on Cancer-Myth to $80\%$, but at the cost of misidentifying presuppositions in $41\%$ of Cancer-Myth-NFP questions and causing a $10\%$ relative performance drop on other medical benchmarks. These findings highlight a critical gap in the reliability of LLMs, show that prompting alone is not a reliable remedy for false presuppositions, and underscore the need for more robust safeguards in medical AI systems.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive performance on various medical benchmarks. Today’s benchmarks primarily focus on medical exam questions or consumer-searched queries, such as MedQA (Jin et al., 2020) and HealthSearchQA (Singhal et al., 2023). Meanwhile, LLMs are increasingly utilized as real-life medical advisors. A recent survey shows that 32.6% of patients now ask LLMs for advice (Mendel et al., 2025), particularly for severe diseases like cancer, where medical resources are limited (Bourgeois et al., 2024; Belong.Life, 2024). However, real-world patient inquiries differ significantly from traditional medical benchmarks, as they often contain patient details and sometimes false presuppositions (Kaplan, 1978), i.e., misconceptions that a patient believes and assumes when asking a question. To effectively respond to real-world patient inquiries, an LLM should (1) provide factual and helpful answers and (2) identify and clarify any misconceptions in the patient’s question. Both aspects are essential for ensuring patient safety and effective medical communication.
@@
To systematically study this issue, we compile a collection of 994 common cancer myths to create an adversarial Cancer-Myth dataset of 585 examples designed to evaluate LLM and medical agent performance in handling patient questions with embedded misconceptions. During this process, we also collect a 150-question Cancer-Myth-NFP set, which LLMs flagged as containing false presuppositions but were confirmed by expert physicians to have none. We initialize the adversarial datasets with a few failure examples from our previous CancerCare study. Using an LLM generator, we create patient questions for each myth, integrating false presuppositions with complex patient details to challenge the models. The LLM responder answers these questions, while a verifier evaluates the response’s ability to address false presuppositions effectively. Responses that fail to correct the presuppositions are added to the adversarial set (i.e., candidates for Cancer-Myth), while successful ones are placed in the non-adversarial set, for use in subsequent generator prompting rounds. We perform three separate runs over the entire set of myths, each targeting GPT-4o, Gemini-1.5-Pro, and Claude-3.5-Sonnet, respectively. The generated questions are finally reviewed by physicians to ensure their relevance and reliability, and are categorized into 7 categories for error analysis by question type.
Our experimental results reveal a lack of awareness in identifying and correcting patient false presuppositions during LLMs’ medical reasoning. No model can fully correct more than 43% of the false presuppositions, even advanced medical agentic methods do not prevent LLMs from ignoring false presuppositions. Typical mitigation strategies, such as adding precautionary prompts with GEPA optimization (Agrawal et al., 2025), can improve the identification rate of false presupposition on Cancer-Myth to 80% with Gemini-2.5-Pro, but they also trigger a 41% rate of incorrect presupposition identification on Cancer-Myth-NFP, which clinicians marked as containing no false presuppositions. Additionally, these strategies induce an average 10% relative performance drop across multiple standard medical benchmarks, including MedQA (Jin et al., 2020), PubMedQA (Jin et al., 2019), SymCat (Al-Ars et al., 2023), Medbullets (Chen et al., 2024), and Crafted-MD (Chen et al., 2024). This suggests that prompting alone is not a straightforward fix for false presuppositions in medical LLMs. In particular, models perform worst on the Inevitable Side Effect category questions, where patient misconceptions, such as assuming a specific treatment will inevitably cause a certain side effect, often go unchallenged by LLMs. These findings underscore the need for improved LLM training and evaluation methods that emphasize patient-centered communication and misinformation detection.
@@
2 Physician Evaluation of LLM Responses to Patient Questions
@@ 2.1 CancerCare data preparation
We selected 25 representative oncology questions sourced from CancerCare website, [^1] https://www.cancercare.org/questions which provides support and resources to individuals affected by cancer (See Appendix A). The selected questions focus on treatment advice and side effects, thus requiring medical expertise. We confirmed these questions cannot be answered by a simple Google web search. We asked three hematology oncology physicians to evaluate four responses to each question: three responses from frontier LLMs GPT-4-Turbo (OpenAI, 2023), Gemini-1.5-Pro (Google, 2023), LLaMa-3.1-405B (Meta, 2024b) and the human responses from the website for comparison. The human answers are provided by licensed medical social workers, who excel in offering compassionate support and general guidance but may lack the depth of medical knowledge for complex oncological queries. Their responses average 237 words in length. To ensure the physicians remained blind to the response source, we prompted LLMs to provide responses of similar length and removed identifiers such as “As an AI chatbot”.
@@ 3 Cancer-Myth Dataset Creation
We develop a dataset to systematically evaluate LLM performance in answering questions with false presuppositions. We collect 994 cancer-related myths and use LLMs to generate adversarial examples based on these myths. These examples are then verified by hematology oncology physicians to ensure validity.
@@ 3.1 Myth collection
To construct a comprehensive collection of cancer-related myths, we first compile a list of 127 cancer types from the official National Cancer Institute (NCI) website [^2] https://www.cancer.gov/types. For each type, we conduct online searches to identify common myths specifically related to the treatment of that type of cancer. All information is collected from publicly accessible online resources. For example:
@@
Source: https://healthmatters.nyp.org/seven-myths-about-colorectal-cancer/.
@@
Figure 4: Example question and information to correct the false presuppositions per category. The proportion of each category in Cancer-Myth is indicated in parentheses.
@@ 3.3 Categorization and expert validation of dataset
To effectively analyze the types of false presuppositions in patient questions, we manually reviewed a subset of 76 examples from our adversarial set, identifying six major categories of false presuppositions: Only/Standard Treatment, No Treatment, Inevitable Side Effect, Causal Misattribution, Underestimated Risk, and No Symptom, No Disease. Examples not fitting these categories are classified as Others. We prompted GPT-4o to categorize each example, achieving an agreement rate of $\sim 90\%$ on the manually annotated subset. Therefore, GPT-4o is used to categorize the remaining examples. Figure 4 lists the questions and the corresponding information to correct the false presuppositions in each category.
To prevent any category from dominating the dataset, we employ the algorithm described in Appendix B to balance the categories before they are verified physicians, ensuring a diverse and representative benchmark. Finally, hematology oncology physicians review adversarial examples to confirm that the following criteria are satisfied: (1) the generated patient question contain a false presupposition; (2) the factual correction is medically sound and appropriately address the false presupposition. Note that criterion (1) is a prerequisite for the assessment of criterion (2). Examples meeting both criteria constitute the Cancer-Myth dataset, whereas those satisfying criterion (1) but failing criterion (2) are discarded. Examples failing criterion (1) form the Cancer-Myth-NFP dataset. We performed an inter-physician agreement analysis on a small batch of 90 examples, which shows 83% of examples had full agreement across all three physicians, and Gwet’s AC1 was 0.78, indicating substantial agreement (Wongpakaran et al., 2013).
To evaluate the realism of Cancer-Myth questions, we conducted a survey comparing it with real CancerCare questions. We randomly selected 10 questions from each source, paired them, and asked 10 NLP researchers to judge which question in each pair appeared more human-written. On average, participants selected the real human-written question 67% of the time. Notably, only 2 out of 10 pairs had over 80% agreement on the human-written question ($p<0.05$). In contrast, for 6 pairs, at least 40% of participants were unable to reliably distinguish the source ($p>0.37$), suggesting that our Cancer-Myth questions resemble real patient questions from CancerCare reasonably well.
Figure 5: (a) GPT-5 performs the best but no frontier LLM corrects the false presuppositions in the patient question more than 43% of the time; Multi-agent medical collaboration does not prevent LLMs from ignoring false presuppositions. (b) Adversarial data generated by Gemini-1.5-Pro causes failures in GPT-4o, but data generated by GPT-4o affects Gemini-1.5-Pro less.
@@ 4.1 Dataset statistics and quality
We generated a total of 1,692 adversarial examples: 868 from GPT-4o, 393 from Gemini-1.5-Pro, 429 from Claude-3.5-Sonnet and 2 manually crafted (i.e., the initial set). After categorization and label balancing, we selected 888 examples for verification by hematology oncology physicians. They filtered these down to 585 valid data points for Cancer-Myth, and 150 no-false-presupposition data points for Cancer-Myth-NFP (153 discarded). We list the final distribution of question categories, generator sources and qualitative examples in Appendix B.
@@ 4.2 Models and metrics
We use GPT-4o as the verifier to evaluate the response of 6 model families: GPT (OpenAI, 2024; 2023), Claude (Anthropic, 2024), DeepSeek (DeepSeek-AI, 2025), Gemini (Google, 2023; 2024; 2025), LLaMA (Meta, 2024a) and Qwen (Qwen, 2024), in total 17 models. We follow the same scoring system as in generation (§3.2) and output a score $s$ in $\{1,0,-1\}$. Additionally, we evaluate an adaptive multi-agent framework MDAgents (Kim et al., 2024), which dynamically assigns collaboration structures among LLMs based on the complexity of the medical task, and achieves state-of-the-art performance on 7 medical benchmarks. Since we expect LLMs to address patient misconceptions without any in-context examples in real patient interactions, all models are evaluated in a zero-shot setup.
Our primary evaluation metric, the Presupposition Correction Score (PCS), is the average score from the verifier. To better align with human (Appendix C) and reduce the difficulty of identifying no correction (s=1s=-1) and partial correction ($s=0$), we introduce the Presupposition Correction Rate (PCR), which focuses solely on fully correct scenarios.
@@
Figure 6: Category-wise Presupposition Correction Score for each model. All models perform poorly on No Treatment and Inevitable Side Effect.
@@ 4.3 Evaluation results
We observe no frontier LLM corrected more than 43% of false presuppositions in patient questions (Figure 5(a)). Among all models, GPT-5 achieves the highest Presupposition Correction Rate (PCR) at 42.1%, followed by Gemini-2.5-Pro at 41.4%, Claude-4-Sonnet at 40.0% and Gemini-1.5-Pro at 27.2%. Surprisingly, GPT-4o shows a low performance on Cancer-Myth, with a PCR of only 5.8%. While GPT-4-Turbo is the best model for providing factual and helpful medical responses (Figure 2), these results indicate that LLM ability to correct false presuppositions does not align with medical knowledge understanding and utilization.
@@
In our experiment, the use of the MDAgents (Kim et al., 2024) does not improve performance. Responses generated through GPT-4o-based MDAgents performed worse than standalone GPT-4o outputs, suggesting that agent-based orchestration is insufficient to address embedded misconceptions.
@@
While overall performance remains limited, GPT-5 achieves the highest PCS among the evaluated models. Its performance advantage primarily comes from better handling of the latter three categories—Causal Misattribution, Underestimated Risk, and No Symptom, No Disease—where it demonstrates greater capacity to detect and explain misleading assumptions. These categories often involve more subtle or technical misconceptions, such as confusing symptom origins or believing that absence of symptoms negates the need for screening. Nonetheless, even in these areas, the best model still leaves substantial room for improvement.
Table 1: GEPA or agent-based precaution improves Cancer-Myth accuracy but reduces performance on other benchmarks, particularly Cancer-Myth-NFP.
Model Method Cancer-Myth Cancer-Myth-NFP MedQA PubMedQA SymCat Medbullets Craft-MD
GPT-4o Plain 12 88 70 67 70 68 55
GEPA 68 59 63 59 61 62 46
Gemini-2.5-Pro Plain 41 96 92 82 91 80 68
GEPA 88 68 85 78 87 72 58
GPT-4o w/ Plain 2 90 89 77 91 82 66
MDAgents Monitor 81 35 86 73 89 80 63
4.4 Mitigation Strategy Evaluation
As LLMs should be able to address patient misconceptions without any in-context examples in real patient interactions, we evaluate on two mitigation strategies: (1) precautious statement-based prompt optimization with GEPA (Agrawal et al., 2025) on a mix of 7 medical benchmarks, Cancer-Myth, Cancer-Myth-NFP, MedQA, PubMedQA, SymCat, Medbullets, and Craft-MD. We take 5 examples per dataset for training and 5 examples for validation. (2) adding a monitoring precautious agent in the MDAgents framework. More details are in Appendix C.
Table 1 shows that precautious statement optimization can improve the performance on Cancer-Myth to 80% with Gemini-2.5-Pro but induce a $28\%$ performance drop in Cancer-Myth-NFP and 55-1515% relative performance drops on other medical benchmarks. On the other hand, modifying the agentic workflow incurs less performance change on standard medical benchmarks, but is drastically overcautious, i.e., identifying most (65%) of the questions with no false presuppositions as containing false presuppositions.
@@ 5 Related Works
LLM sycophancy.
Questions containing false presuppositions have long been studied in linguistic literature (Kaplan, 1978), where the appropriate and unambiguous response is a corrective response that negates the false presupposition. LLMs, however, exhibit sycophancy—a tendency to align with users’ opinions (Perez et al., 2023)—making them prone to accepting and reinforcing false presuppositions in questions. Common sycophancy mitigation strategies, such as in-context learning (Zhao, 2023), precautionary prompts (Varshney et al., 2024), and augmenting contextual knowledge from LLMs (Luo et al., 2023), are ineffective in zero-shot settings (Rrv et al., 2024), which are crucial for real-world patient-LLM interactions. Recently, Yu et al. (2023) introduced Crepe, an open-domain QA benchmark that targets false presuppositions in Reddit questions. Building on this idea, subsequent benchmarks have examined model robustness to presuppositional errors in different settings. For example, $(\textsc{QA})^{2}$ (Kim et al., 2023) focuses on frequently searched queries, while Vu et al. (2024) proposed FreshQA, a dynamic benchmark featuring factually incorrect presuppositions that require explicit rebuttal. In the medical domain, Srikanth et al. (2024) introduced Pregnant Questions, where mothers ask about pregnancy and infant care. To our knowledge, Cancer-Myth is the first benchmark specifically designed to evaluate how LLMs handle false presuppositions in the cancer care domain.
Adversarial generation.
Semantic, manual, or rule-based adversarial data generation (Jia & Liang, 2017; Rajpurkar et al., 2018) predates the emergence of LLMs, yet perturbation methods remain effective for creating challenging negatives (Zhu et al., 2022; Fu et al., 2025). Recently, automatic LLM-based synthetic adversarial generation (Bartolo et al., 2021; Fu et al., 2023) and filtering (Bras et al., 2020) have enhanced model robustness against attacks and mitigated overfitting. Dynabench (Kiela et al., 2021) combined human annotation with model inspection to generate misleading examples. Moreover,  Sung et al. (2024a; b) underscored the necessity of human involvement and proposed criteria to ensure that adversarialness appropriately targets models rather than humans. To balance labor efficiency with robustness, we adopt a hybrid strategy: leveraging an LLM verifier in conjunction with physician annotation to filter questions containing false presuppositions.